OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –2x2 – 32x – 126?

OpenStudy (anonymous):

plzz <3

OpenStudy (anonymous):

The vertex is (-8,2) (using the methods from your previos post) and the parabola opens downward: Range is \[(-\infty,2]\]Domain is \[(-\infty, \infty)\]

OpenStudy (anonymous):

is it less than or equal to 2 or greater than or equl to 2 andthank u u r so sweet :)

OpenStudy (anonymous):

less than or equal to 2

jimthompson5910 (jim_thompson5910):

Range can also be written as \[\Large \{y|y\le2\}\]

OpenStudy (anonymous):

you are a genius